OpenStudy (anonymous):

Functions- If 2f(x) - 3f(1/x) = x^2, determine f(2)

OpenStudy (jamesj):

First write a = f(2) and b = f(1/2) Setting first x = 2, from your equation: 2f(2) - 3f(1/2) = 4 i.e., 2a - 3b = 4 Set now x = 1/2, 2f(1/2) - 3f(2) = 1/4 i.e., 2b - 3a = 1/4 You now have two equations in two unknowns, a and b. Solve. Got it?